geerky42 (geerky42):

\[\Huge 1 = 2 \text{ via Continued Fractions}\]

geerky42 (geerky42):

This is true of course that \(1=\dfrac{2}{3-1}\) Now, let's substitute this very expression for 1 in the denominator: \[1=\dfrac{2}{3-\dfrac{2}{3-1}}\] We can do that one more time: \[1 = \dfrac{2}{3-\dfrac{2}{3-\dfrac{2}{3-1}}}\] And one more time to make sure there is no misunderstanding of the construction, \[1 = \dfrac{2}{3-\dfrac{2}{3-\dfrac{2}{3-\dfrac{2}{3-1}}}}\] At this point I am assuming that further steps could be performed by any reader who got this far. To indicate that possibility I'll use the ellipsis: \[1 = \dfrac{2}{3-\dfrac{2}{3-\dfrac{2}{3-\dfrac{2}{\ddots}}}}\] Well, we also know that \(2 =\dfrac{2}{3-2}\). With this as a starting point, we follow in the footsteps of the previous example. Replacing 2 in the denominator with that expression gives \(2=\dfrac{2}{3-\dfrac{2}{3-2}}\) To continue as before: \(2=\dfrac{2}{3-\dfrac{2}{3-\dfrac{2}{3-2}}}\) And finally, \(2=\dfrac{2}{3-\dfrac{2}{3-\dfrac{2}{3- \dfrac{2}{\ddots } }}}\) But this is exactly the same continued fraction. By necessity we conclude that 1=2.

OpenStudy (openstudier):

We can think of this as: \[1=\frac{ 2 }{ 3-\frac{2 }{ x } }\] And, we still get 1=2.

geerky42 (geerky42):

True.

OpenStudy (openstudier):

Oh my god, we cheated math.

geerky42 (geerky42):

I await someone to disprove this.

geerky42 (geerky42):

@Hero @dumbcow @robtobey @blues @ajprincess @Yahoo! @AccessDenied

OpenStudy (anonymous):

Goddamn, 1 is NOT equal 2. PERIOD.

geerky42 (geerky42):

Well, can you disprove it?

OpenStudy (anonymous):

It's not the same continued fraction... The scheme is different... Are you happy now?

geerky42 (geerky42):

Any thoughts, @LogicalApple ?

OpenStudy (anonymous):

I know x = 2/(3 - x) yields 1 and 2 as solutions. I have never worked with continuous fractions before but the little bit of research tells me that a rational number has a finite continued fraction and an irrational one has an infinite continued fraction. Although what you have displayed does not disprove either fact.

OpenStudy (anonymous):

Maybe continuous fractions are not unique?

OpenStudy (anonymous):

These are generalized continued fractions, by the way. In continuous fractions, every numerator has to be 1.

OpenStudy (anonymous):

Also, an infinite continued fraction represents by definition an irrational number, which 1 and 2 are not.

OpenStudy (anonymous):

also, the continued fraction represents the decimal part of the number, which in the case of 1 and 2 should be 0

OpenStudy (anonymous):

you're redefining continued fraction.

geerky42 (geerky42):

You basically said the same thing as @micahwood50 did. Thank for post, though.

OpenStudy (dumbcow):

no matter how far you continue the recursive substitution, at the very bottom they will be unique 3-1 not equal 3-2 and to infer infinite regression implies an irrational number as was posted earlier, thus your conclusion is false since you can't equate an integer to an irrational number plus, any 4 yr old will tell you 1 of something is not the same as 2 of something :)